# Evaluating Numerical Reasoning in Text-to-Image Models

### v1

@@ B.1.3 Prompt structure: Additive prompts
| Dalle 3 | Imagen-C | Muse-B | |||||||
| Ground-truth number | 1 | 2 | 3 | 1 | 2 | 3 | 1 | 2 | 3 |
| numeric-simple2-additive | -7.3 | -41.3 | -40.3 | -30.9 | -46.8 | -23.6 | -2.2 | -22.4 | -10.3 |
| 2-additive3-additive | -7.3 | -9.8 | -4.6 | 1.3 | -6.6 | -14.8 | -11.4 | -7.8 | -7.8 |
| numeric-simple3-additive | -14.5 | -51.1 | -45.0 | -29.6 | -53.4 | -38.4 | -13.6 | -30.1 | -18.1 |
Figure A4 shows drops in accuracy when a specific number (indicated on the x-axis) occurs in different prompt types. In addition, Table A5 shows the exact drops in accuracy for selected models for the numbers: “1”, “2” and “3”. While for some models the drop in accuracy for the number “1” (c.f. the three bars above the x-tick “1”) is much smaller (e.g. DALL·E 3) or non-significant (Muse-B) compared to accuracy drops for some other models (e.g. Imagen-A, Imagen-C), all models show a substantial and significant drop when “2” is present in the prompt (c.f. the three bars above the x-tick “2”).
@@ C.2 Quantitative Analysis
In total, in Annotation Task 1 we collected annotations, ranging from approximately K annotations per model, except for Imagen-C where we collected K annotations as five additional images per prompt were generated for evaluation and fine-tuning experiments in Section 6 (for a total of 10 seeds for Imagen-C). In Annotations Tasks 2 and 3, we collected and annotations, respectively. The analyses and statistics below will be based on using five seeds for each model, although the data we release also includes additional images and annotations. As well, released data will not contain any images showing faces (approximately 122 images), due to potential privacy concerns, but analyses below were done including annotations collected for those images.
@@
We observe high level of agreement among annotators. In Task 1, for each image–question pair, all 5/5 raters gave the same response in 79% of cases, at least 4/5 raters gave the same response in 89% of cases and at least 3/5 in 97%. In Task 2, the analogous percentages were as follows: 79%, 92% and 99%, and in Task 3 where we only had binary responses all 5/5 raters agreed in 85% of cases, while at least 4/5 agreed in 95%. Expressed as Krippendorff’s alpha coefficient of inter-annotator agreement, for Task1, for Task 2, and for Task 3. A breakdown per model annotator agreements in Task 1 is shown in Table A10.
@@ D.2.2 Fine-tuning PaLIGemma on GeckoNum
![[Uncaptioned image]](extracted/5682110/figs/appendix/results/fail_spoon.png)
![[Uncaptioned image]](extracted/5682110/figs/appendix/results/unin_pencil.jpg)
![[Uncaptioned image]](extracted/5682110/figs/appendix/results/over_okra.png)
![[Uncaptioned image]](extracted/5682110/figs/appendix/results/1_cinammon_stick_159_2.png)
![[Uncaptioned image]](extracted/5682110/figs/appendix/results/back_samosa.png)
![[Uncaptioned image]](extracted/5682110/figs/appendix/results/fail_corckscrew.png)
![[Uncaptioned image]](extracted/5682110/figs/appendix/results/unin_corckscrew.png)
![[Uncaptioned image]](extracted/5682110/figs/appendix/results/over_chocolate.png)
![[Uncaptioned image]](extracted/5682110/figs/appendix/results/one_manatee_9_0.png)
![[Uncaptioned image]](extracted/5682110/figs/appendix/results/one_kangaroo.png)

### v2

@@ Acknowledgments.
Checklist
1.
For all authors…
Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
Did you describe the limitations of your work? [Yes]
Did you discuss any potential negative societal impacts of your work? [N/A]
Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
2.
If you are including theoretical results…
Did you state the full set of assumptions of all theoretical results? [N/A]
Did you include complete proofs of all theoretical results? [N/A]
3.
If you ran experiments (e.g. for benchmarks)…
Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] https://github.com/google-deepmind/geckonum_benchmark_t2i
Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] Section 6
Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes]
Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes]
4.
If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
If your work uses existing assets, did you cite the creators? [Yes]
Did you mention the license of the assets? [Yes]
Did you include any new assets either in the supplemental material or as a URL? [Yes]
Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes]
Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes]
5.
If you used crowdsourcing or conducted research with human subjects…
Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] See Appendix C.1
Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [Yes]
Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [Yes]
[appendices]
@@ B.1.3 Prompt structure: Additive prompts
| SN2A | 2A3A | SN3A | ||
| Number | ||||
| DALLE 3 | 1 | -9.7 | -3.2 | -12.9 |
| 2 | -19.2 | -10.4 | -29.6 | |
| 3 | -14.2 | -9.5 | -23.7 | |
| Imagen-C | 1 | -17.4 | -5.6 | -23.0 |
| 2 | -35.4 | -0.1 | -35.5 | |
| 3 | -21.8 | -9.5 | -31.4 | |
| Midjourney v6 | 1 | -27.9 | 9.3 | -18.6 |
| 2 | -16.7 | -13.0 | -29.7 | |
| 3 | -17.9 | -10.8 | -28.6 | |
| Muse-B | 1 | -0.1 | -10.3 | -10.4 |
| 2 | -20.7 | 4.2 | -16.5 | |
| 3 | -1.6 | -4.1 | -5.7 | |
| SD3 | 1 | -4.7 | -1.7 | -6.4 |
| 2 | -32.8 | -5.0 | -37.8 | |
| 3 | -18.5 | -6.0 | -24.6 |
Figure A4 shows drops in accuracy when a specific number occurs in different prompt types. In addition, Table A5 shows the exact drops in accuracy for selected models for the numbers: “1”, “2” and “3”. While for some models the drop in accuracy for the number “1” (c.f. the three bars above the x-tick “1”) is much smaller (e.g. DALL·E 3) or non-significant (Muse-B) compared to accuracy drops for some other models (e.g. Imagen-A, Imagen-C), all models show a substantial and significant drop when “2” is present in the prompt (c.f. the three bars above the x-tick “2”).
@@ C.2 Quantitative Analysis
| SD1.5 | 0.2 | 4.2 | 10.1 | 11.5 | 73.9 |
| SD2.1 | 0.1 | 4.5 | 11.8 | 12.6 | 71.0 |
| SDXL | 0.3 | 5.1 | 9.7 | 11.0 | 73.9 |
| SD3 | 0.2 | 3.8 | 9.7 | 11.3 | 75.0 |
In total, in Annotation Task 1 we collected annotations, ranging from approximately K annotations per model, except for Imagen-C where we collected K annotations as five additional images per prompt were generated for evaluation and fine-tuning experiments in Section 6 (for a total of 10 seeds for Imagen-C). In Annotations Tasks 2 and 3, we collected and annotations, respectively. For Imagen-C, the analyses and statistics below will be based on using five seeds for each model, although the data we release also includes additional images and annotations. Released data will not contain any images showing faces (approximately 122 images), due to potential privacy concerns, but analyses below were done including annotations collected for those images.
@@
We observe high level of agreement among annotators. In Task 1, for each image–question pair, all 5/5 raters gave the same response in 77.5% of cases, at least 4/5 raters gave the same response in 87.9% of cases and at least 3/5 in 96.5%. In Task 2, the analogous percentages were as follows: 74.7%, 89.5% and 99.3%, and in Task 3 where we only had binary responses all 5/5 raters agreed in 84.9% of cases, while at least 4/5 agreed in 94.9%. Expressed as Krippendorff’s alpha coefficient of inter-annotator agreement, for Task1, for Task 2, and for Task 3. A breakdown per model annotator agreements in Task 1 is shown in Table A10.
@@ D.2.2 Fine-tuning PaLIGemma on GeckoNum
![[Uncaptioned image]](extracted/6048850/figs/appendix/results/fail_spoon.png)
![[Uncaptioned image]](extracted/6048850/figs/appendix/results/unin_pencil.jpg)
![[Uncaptioned image]](extracted/6048850/figs/appendix/results/over_okra.png)
![[Uncaptioned image]](extracted/6048850/figs/appendix/results/1_cinammon_stick_159_2.png)
![[Uncaptioned image]](extracted/6048850/figs/appendix/results/back_samosa.png)
![[Uncaptioned image]](extracted/6048850/figs/appendix/results/fail_corckscrew.png)
![[Uncaptioned image]](extracted/6048850/figs/appendix/results/unin_corckscrew.png)
![[Uncaptioned image]](extracted/6048850/figs/appendix/results/over_chocolate.png)
![[Uncaptioned image]](extracted/6048850/figs/appendix/results/one_manatee_9_0.png)
![[Uncaptioned image]](extracted/6048850/figs/appendix/results/one_kangaroo.png)